OpenStudy (anonymous):

I need some help please solve the inequality. graph the solution on a number line and check your results −7x≥63

OpenStudy (anonymous):

yeah your back

OpenStudy (jiteshmeghwal9):

\[-7x \ge 63\]\[x \ge 63/-7\]\[x \ge -9\]

OpenStudy (jiteshmeghwal9):